OpenStudy (anonymous):

find the value for x and y for which (x,2) and (8,y) are points on the graph \[-4x-3y=3\]

OpenStudy (precal):

sub y=2 and solve for x sub x = 8 and solve for y you can do it

OpenStudy (precal):

gotta go

OpenStudy (anonymous):

x=-4/9 and y=-35/3 ??

OpenStudy (anonymous):

http://cs.selu.edu/~rbyrd/math/intercept/

OpenStudy (anonymous):

it should be x=-9/4 .

OpenStudy (anonymous):

yea i meant -9/4

OpenStudy (anonymous):

thanks!

OpenStudy (precal):